The lenght of a retangle is 3 cm longer than twice the width.the area of the retangle is 90 cm ^2
Answer by user_dude2008(1862)

"The lenght of a retangle is 3 cm longer than twice the width" --> L=2W+3

A=LW=(2W+3)(W)=90

2W^2+3W=90

2W^2+3W-90=0

(W-6)(2W+15)=0

W=6 or W=-15/2 ---> L=2(6)+3=15

Answer: width = 6, length = 15